A ball is thrown making an angle of 30 degrees above the horizontal with a velocity of 20m/s from the top of a building 50m high. Calculate



(i) The time of flight (ii) The horizontal range. (iii) The velocity with which it hits the ground. (g= 10m/s²)

1
Expert's answer
2022-01-31T10:13:58-0500

(g=9.8 m/s²)

(i) The time of flight:


"t_{up}=\dfrac{v\sin\theta}{g},\space t_{down}=\sqrt{\dfrac{2[h+v^2\sin^2\theta/(2g)]}{g}},\\\space\\
T=t_{up}+t_{down}=4.05\text{ s}."

(ii) The horizontal range:


"R=v\cos\theta·T=70.2\text{ m}."

(iii) The velocity with which it hits the ground.


"u=\sqrt{(v\cos\theta)^2+2g[h+v^2\sin^2\theta/(2g)]}=\\=37\text{ m/s}^2."